# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v1

@@ Abstract
In federated learning, the heterogeneity of client data has a great impact on the performance of model training. Many heterogeneity issues in this process are raised by non-independently and identically distributed (Non-IID) data. This study focuses on the issue of label distribution skew. To address it, we propose a hybrid federated learning framework called HFLDD, which integrates dataset distillation to generate approximately independent and equally distributed (IID) data, thereby improving the performance of model training. Particularly, we partition the clients into heterogeneous clusters, where the data labels among different clients within a cluster are unbalanced while the data labels among different clusters are balanced. The cluster headers collect distilled data from the corresponding cluster members, and conduct model training in collaboration with the server. This training process is like traditional federated learning on IID data, and hence effectively alleviates the impact of Non-IID data on model training. Furthermore, we compare our proposed method with typical baseline methods on public datasets. Experimental results demonstrate that when the data labels are severely imbalanced, the proposed HFLDD outperforms the baseline methods in terms of both test accuracy and communication cost.
@@ I Introduction
In recent years, the rapid advancement of deep neural networks has sparked a revolution across various areas, such as image classification [1, 2], object detection [3, 4] and semantic segmentation [5]. Traditional centralized machine learning methods typically aggregate all data to a central server for model training. However, these methods may involve transmission of a large amount of sensitive data, which can potentially induce a risk of privacy exposure to users. With the development of edge computing and the increasing importance of data privacy protection, federated learning (FL) has become an emerging machine learning paradigm and attracted widespread attentions [6, 7, 8].
@@
In the widely used FL-based method FedAvg [9], the server updates the global model by aggregating the local model parameters with simply weighted averaging. This method performs well when dealing with independently and identically distributed (IID) data. However, researchers have found that data among clients are often Non-IID in practice. In such cases, if the server simply aggregates the local model parameters with weighted averaging, local model drift may make the global model deviate from the global optimum [10], leading to a slower convergence of the global model.
@@
To mitigate the impact of Non-IID data on FL-based model training, researchers have explored various approaches. At the Algorithm level, efforts focus on the improvements of loss functions [11, 12, 13] and aggregation methods [14, 15, 16]. However, they cannot fully alleviate the impact of Non-IID data on model training and may have a poor performance when dealing with high statistical heterogeneity. At the system level, client clustering techniques are introduced to create a multi-center framework [17, 18, 19]. These techniques aim to group clients with similar data distributions together to form clusters. Nevertheless, this approach primarily focuses on client similarity and fails to fully exploit the potential advantages of client heterogeneity. At the data level, data sharing is considered an effective strategy [20, 21, 22]. It mitigates the negative impact of Non-IID data by sharing a small amount of local data. Although this method can improve the performance of the global model, the server typically lacks comprehensive knowledge of the clients’ data distributions, making it difficult to obtain a uniformly distributed dataset. Moreover, this approach conflicts with the privacy protection principles of FL.
@@
Recent works on dataset distillation shows promising potential to overcome restrictions on data sharing while protecting privacy. Dataset distillation[23, 24, 25, 26] aims to achieve good generalization performance by training models on a synthesized dataset with much smaller size than the real dataset. Ref. [27] indicates that the privacy of transmitting distilled data should be no worse than transmitting model parameters. Additionally, distilled data not only provides visual privacy but also enhances the model’s resistance to membership inference attacks [28].
@@
Motivated by the above observations, we propose a new learning framework named HFLDD, integrating dataset distillation to hybrid federated learning (HFL). The key chanllenge of our research is how to effectively transmit distilled data among clients to compensate data heterogeneity. Specifically, The main idea of the design is to partition clients into a few clusters where data labels are heterogeneous within each cluster but balanced across different clusters. Furthermore, within each cluster, cluster members transmit their distilled data to a cluster header. In addition, the adoption of client clustering and dataset distillation in FL can reduce the total communication cost of model training, while enhancing the generalization capabilities and performance of model in diverse data environments.
The contributions of our work are as follows:
@@
    We extend the calculation method for communication costs from traditional FL to our proposed HFLDD, and provide a closed-form expression about the communication cost of HFLDD.
@@
The remainder of this paper is organized as follows. Section II introduces related work. Section III provides the preliminaries and problem statement. Section IV introduces the proposed framework HFLDD and Section V shows the performance evaluation. Section VI provides a discussion, and Section VII presents the conclusion.
@@ II-A Federated Learning on Non-IID data
The basic frameworks of FL include traditional FL, decentralized FL (DFL), and HFL. Addressing the performance decline of FL on Non-IID data is a critical challenge. Each framework employs different methods to tackle this issue.
@@
In traditional FL, all the clients train models locally and then transmit these local models to a central server for global aggregation. Various methods have been proposed since the introduction of FedAvg [9] to address this heterogeneity issue. One main direction attempts to improve the design of local loss function. FedProx [12] introduces an additional $L_{2}$ regularization term in the loss function to limit the size of local updates. SCAFFOLD [11] applies the variance reduction technique to correct the local training. MOON [13] utilizes the similarity between model representations to correct the local training of clients. However, when the data heterogeneity is severe, the achieved accuracy remains low for regularization-based and weight adjustment methods.
@@
DFL can implement distributed learning from data distributed across multiple clients without the need for a central server. To address the data heterogeneity issues in DFL, the existing researches mainly focused on how to utilize the topological structure to improve the learning process. D-Cliques[29] addresses label shift problems in decentralized settings by constructing heterogeneous topological structures based on the underlying data distribution of each client, which will induce privacy leakage. In consideration of privacy preservation, Ref. [30] applies a proxy-based method to learn the local datasets’ label knowledge based on a global dataset and construct heterogeneous topological structure for DFL. The proposed method has shown that when the neighbours of each client are with heterogeneous data, it can improve the convergence of model training in addressing label distribution skew problems. This insight inspires the main idea of our method. It should be noted that the paradigm of fully distributed topological structure construction and learning suffers from heavy communication overhead.
@@
Alternative to traditional and decentralized FL, HFL involves both client-to-server communication and client-to-client communication. In [31], a multi-stage HFL algorithm is developed to handle the issue of Non-IID data by continuously adjusting the rounds of client-to-client communication and global aggregation cycles. Particularly, Ref. [32] proposes an HFL framework, called FedSeq, based on client clustering and sequential training. It partitions clients into multiple clusters, with each cluster head communicating with the server. Within each cluster, clients form a ring topology for sequential training. The core idea of FedSeq is that the model in the cluster head is based on the training data from all the clients within this cluster, thereby improving the test accuracy. FedSeq requires multiple rounds of parameter transmission among clients, which still incurs significant communication cost.
@@ II-B Dataset Distillation
The current dataset distillation frameworks are mainly divided into meta-learning frameworks and data matching frameworks [33]. Ref. [23] proposes a meta-learning framework, which adopts a dual optimization approach. However, this method involves a nested loop that can be difficult to optimize. Considering the existence of closed-form solutions in the kernel regression paradigm, Ref. [34] replaces the neural network in the inner loop of the meta-learning framework with a kernel model. In the data matching framework, DC [35, 36] and DM [37, 38] aim to match the gradients and distribution of distilled data with original data during training. Trajectory matching [39] generates synthetic data by matching the training trajectories.
Recently, the integration of dataset distillation into FL has attracted widespread attentions. In DynaFed[40], the server generates distilled data through trajectory matching after each aggregation, and further updates the global model based on this data. Ref. [41] improves communication efficiency and test accuracy by constructing distilled data on each client to locally match the loss landscape of the original data. However, these works do not consider leveraging D2D (device-to-device) networks to compensate for client heterogeneity. Ref. [42] introduces a method for extracting meta knowledge from clients and exchanging it among clients, but this knowledge sharing may be less effective in scenarios with severe data heterogeneity.
In comparison with the aforementioned works, we design a new framework of HFL to address the issue of Non-IID data. It extends the idea of constructing heterogeneous topological structure in DFL to HFL, and fully leverage the potential of heterogeneity among clients.
@@ III-A Dataset Distillation
The main idea of dataset distillation is to transfer a large dataset into a small dataset, while preserving the essential characteristics of the original dataset. Suppose the original dataset and the distilled dataset are respectively denoted by $\mathcal{T}=\left\{\left(\mathrm{x}_{t,i},y_{t,i}\right)\right\}_{i=1}^{|\mathcal{T}|}$ and $\mathcal{S}=\left\{\left(\mathrm{x}_{s,j},y_{s,j}\right)\right\}_{j=1}^{|\mathcal{S}|}$, where $\mathrm{x}_{t,i}\in\mathbb{R}^{d}$ and $y_{t,i}\in\mathbb{R}^{n_{c}}$ are respectively the input data and the corresponding label of the $i$-th sample in the original dataset. $\mathrm{x}_{s,j}\in\mathbb{R}^{d}$ and $y_{s,j}\in\mathbb{R}^{n_{c}}$ are respectively the input data and the corresponding label of the $j$-th sample in the distilled dataset. $|\mathcal{T}|$ and $|\mathcal{S}|$ represent the sizes of the original dataset and the distilled dataset, respectively. Generally, the size of $\mathcal{T}$ is considerably larger than $\mathcal{S}$, i.e.,$|\mathcal{S}|\ll|\mathcal{T}|$.
To realize dataset distillation, one of the most widely used methods is Kernel Inducing Points (KIP) [43, 34], which is a first-order meta-learning algorithm based on Kernel Ridge Regression (KRR). Specifically, KIP method is to find a synthetic dataset $\mathcal{S}$ that is approximate to the target dataset $\mathcal{T}$ by minimizing the following loss function
@@
where $\lambda>0$ is a regularization parameter. $y_{t}\in\mathbb{R}^{|\mathcal{T}|\times n_{c}}$ and $y_{s}\in\mathbb{R}^{|\mathcal{S}|\times n_{c}}$ are the label of $\mathcal{T}$ and $\mathcal{S}$ respectively. ${K}_{X_{s}X_{s}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ and $K_{X_{t}X_{s}}\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{S}|}$ are two kernel matrices defined as $(K_{X_{s}X_{s}})_{ij}=k\left(\mathrm{x}_{s,i},\mathrm{x}_{s,j}\right)$ and $(K_{X_{t}X_{s}})_{ij}=k\left(\mathrm{x}_{t,i},\mathrm{x}_{s,j}\right)$, where $k(\cdot)$ denotes the kernel function. KIP is highly efficient for dataset distillation because it only involves first-order optimization.
@@ III-B Federated Learning
Consider a classic scenario of FL with $N$ participants collaboratively training a global model. Each participant possesses a local dataset, represented as $\mathcal{D}_{i}=\left\{\left(X_{i},y_{i}\right)\right\},i=1,\cdots,N$, where $X_{i}\in\mathbb{R}^{n_{i}\times d}$ and $y_{i}\in\mathbb{R}^{n_{i}\times n_{c}}$ denote the tuple of input data points and their corresponding labels. $n=\sum_{i=1}^{N}{n_{i}}$, where $n_{i}=|\mathcal{D}_{i}|$ is the number of samples in $\mathcal{D}_{i}$. The objective of collaborative learning is to solve the following distributed parallel optimization problem
@@
In the process of training, each participating client trains the local model based on their local dataset and then transmits the local models to a server for global model aggregation. Specifically, in round $t$, client $i$ trains its local model $\omega_{i}^{t}$ by minimizing the local loss function $f_{i}\left(\omega\right)$ using gradient descent method. Subsequently, the server updates the global model $\omega^{t}$ through aggregation rules, e.g., FedAvg. If the client data are IID, the above method can provide the optimal solution of (2). while if the client data are Non-IID, the obtained global model may be drifted from the optimum. The communication cost of the above process comes from the local model uploading and global model downloading. By referring to [27], the communication cost of the widely used FedAvg is
@@
$a_{h}$ cluster header in the $h$-th heterogeneous cluster
$\mathcal{A}$ set of cluster header
$\mathcal{D}_{h}$ whole dataset on the $h$-th cluster header
@@
${\omega}_{h}^{t}$ local model on the $h$-th cluster header at communication round $t$
@@
$E_{1}$ number of epoch in pre-train at label knowledge collection stage
$E_{2}$ number of epoch in local training at model training stage
@@ III-C Problem Statement
In this paper, we consider a typical scenario of distributed learning with Non-IID data, i.e. label distribution skew [44]. Our focus lies in addressing the issue of quantity-based label imbalance, where each client owns data with a fixed number of labels. Supposing $N_{c}$ denotes the total number of label classes, and each client owns data with $C$ different labels. The data labels among different clients could overlap. If $C=N_{c}$, the data labels among all the clients are the same, which can be regarded as quantity-based label balance. if $C<N_{c}$, the data labels among all the clients are imbalanced. If $C=1$, it represents the worst-case scenario of quantity-based label imbalance, where the local dataset on each client contains only a single label. It has been verified that the quantity-based label imbalance will decrease the performance of model training to a great extent [12, 45].
Instead of traditional FL, we will design a new HFL framework that integrates dataset distillation to create approximately label-balanced local datasets, thereby alleviating global model drift caused by Non-IID data. Furthermore, it should be noted that the clients can communicate with each other.

### v2

@@ Abstract
With the development of edge computing, Federated Learning (FL) has emerged as a promising solution for the intelligent Internet of Things (IoT). However, applying FL in mobile edge-cloud networks is greatly challenged by statistical heterogeneity and high communication overhead. To address it, we propose a hybrid federated learning framework called HFLDD, which integrates dataset distillation to generate approximately independent and equally distributed (IID) data, thereby improving the performance of model training. In particular, we partition the clients into heterogeneous clusters, where the data labels among different clients within a cluster are unbalanced while the data labels among different clusters are balanced. The cluster heads collect distilled data from the corresponding cluster members, and conduct model training in collaboration with the server. This training process is like traditional federated learning on IID data, and hence effectively alleviates the impact of non-IID data on model training. We perform a comprehensive analysis of the convergence behavior, communication overhead, and computational complexity of the proposed HFLDD. Extensive experimental results based on multiple public datasets demonstrate that when data labels are severely imbalanced, the proposed HFLDD outperforms the baseline methods in terms of both test accuracy and communication cost.
@@ I Introduction
In recent years, the rapid advancement of IoT devices and smart terminals has led to an exponential increase in global data volume [1]. Mobile edge computing (MEC) provides an efficient solution by offloading data storage, analysis, and processing to the network edge, significantly improving efficiency and response times [2]. Meanwhile, advances in deep neural networks have advanced tasks such as image classification [3, 4], object detection [5, 6], and semantic segmentation [7], making machine learning (ML) workloads a primary focus in MEC. To address challenges such as privacy leakage and high communication costs, Federated Learning (FL) has become an emerging machine learning paradigm and has attracted widespread attention [8, 9, 10].
@@
In the widely used FL-based method FedAvg [11], the server updates the global model by aggregating the local model parameters with simply weighted averaging. This method performs well when dealing with independently and identically distributed (IID) data. However, researchers have found that the data among clients is often non-IID in practice. In such cases, if the server simply aggregates the parameters of the local model with weighted averaging, the drift of the local model may cause the global model to deviate from the global optimum [12], leading to slower convergence of the global model.
@@
To mitigate the impact of non-IID data on FL-based model training, researchers have explored various approaches. At the algorithm level, efforts focus on improving loss functions [13, 14, 15] and aggregation methods [16, 17, 18]. However, they cannot fully alleviate the impact of non-IID data on model training and may have poor performance when dealing with high statistical heterogeneity. At the system level, client clustering techniques are introduced to create a multicenter framework [19, 20, 21]. These techniques aim to group clients with similar data distributions together to form clusters. However, this approach primarily focuses on the similarity of the client and does not fully exploit the potential advantages of the heterogeneity of the client. At the data level, data sharing is considered an effective strategy [22, 23, 24]. It mitigates the negative impact of non-IID data by sharing a small amount of local data. Although this method can improve the performance of the global model, the server typically lacks comprehensive knowledge of the client data distributions, making it difficult to obtain a uniformly distributed dataset. Moreover, this approach conflicts with the privacy protection principles of FL.
@@
Recent work on data collection shows promising potential to overcome restrictions on data sharing while protecting privacy. Dataset distillation[25, 26, 27, 28] aims to achieve good generalization performance by training models on a synthesized dataset much smaller than the real dataset. Ref. [29] indicates that the privacy of the transmission of distilled data should be no worse than the transmission of model parameters. Additionally, distilled data not only provides visual privacy, but also enhances the model’s resistance to membership inference attacks [30].
@@
Motivated by the above observations, we propose a new learning framework named HFLDD, integrating dataset distillation with hybrid federated learning (HFL). The key challenge of our research is how to effectively transmit distilled data among clients to compensate data heterogeneity. Specifically, the main idea of the design is to partition clients into a few clusters where data labels are heterogeneous within each cluster but balanced across different clusters. Furthermore, within each cluster, the cluster members transmit their distilled data to a cluster head. In addition, the adoption of client clustering and dataset distillation in FL can reduce the total communication cost of model training, while enhancing the generalization capabilities and performance of the model in diverse data environments.
The contributions of our work are as follows.
@@
    We theoretically analyze the convergence behavior, communication overhead, and computational complexity of the proposed HFLDD. Specifically, we derive a convergence bound and provide a closed-form expression for the communication cost associated with HFLDD. Additionally, we explicitly characterize the computational complexity for the server, the cluster head, and the cluster member, respectively.
@@
The remainder of this paper is organized as follows. Section II introduces the related work. Section III provides the preliminaries and problem statement. Section IV presents the proposed HFLDD framework in detail. Section V provides theoretical analysis on the convergence, communication cost, and computational complexity of HFLDD. Section VI conducts extensive experimental evaluations to validate the effectiveness of HFLDD. Section VII provides a discussion. Finally, Section VIII presents the conclusion.
@@ II-A Federated Learning on non-IID data
The basic frameworks of FL include traditional FL, decentralized FL (DFL), and HFL. Addressing the performance decline of FL on non-IID data is a critical challenge. Each framework employs different methods to address this problem.
@@
In traditional FL, all clients train models locally and then transmit these local models to a central server for global aggregation. Various methods have been proposed since the introduction of FedAvg [11] to address this heterogeneity issue. One main direction attempts to improve the design of the local loss function. FedProx [14] introduces an additional $L_{2}$ regularization term in the loss function to limit the size of local updates. SCAFFOLD [13] applies the variance reduction technique to correct local training. MOON [15] uses the similarity between model representations to correct local client training. However, when the data heterogeneity is severe, the accuracy achieved remains low for regularization-based and weight-adjustment methods.
@@
DFL can implement distributed learning from data distributed across multiple clients without the need for a central server. To address data heterogeneity issues in DFL, existing research focused primarily on how to utilize the topological structure to improve the learning process. D-Cliques[31] address label shift problems in decentralized settings by constructing heterogeneous topological structures based on the underlying data distribution of each client, which will induce privacy leakage. In consideration of privacy preservation, Ref. [32] applies a proxy-based method to learn the knowledge of local dataset labels based on a global dataset and to construct a heterogeneous topological structure for DFL. The proposed method has shown that when the neighbors of each client have heterogeneous data, it can improve the convergence of model training in addressing label distribution skew problems. This insight inspires the main idea of our method. It should be noted that the paradigm of fully distributed topological structure construction and learning suffers from a high communication overhead.
@@
Unlike traditional and decentralized FL, HFL involves both client-to-server communication and client-to-client communication. In [33], a multistage HFL algorithm is developed to handle the issue of non-IID data by continuously adjusting the rounds of client-to-client communication and global aggregation cycles. In particular, Ref. [34] proposes an HFL framework, called FedSeq, based on client clustering and sequential training. It partitions clients into multiple clusters, with each cluster head communicating with the server. Within each cluster, clients form a ring topology for sequential training. The core idea of FedSeq is that the model in the cluster head is based on the training data of all the clients within this cluster, thereby improving the test accuracy. FedSeq requires multiple rounds of parameter transmission among clients, which still incurs significant communication cost.
@@ II-B Dataset Distillation
The current dataset distillation frameworks are mainly divided into meta-learning frameworks and data matching frameworks [35]. Ref. [25] proposes a meta-learning framework that adopts a dual optimization approach. However, this method involves a nested loop that can be difficult to optimize. Considering the existence of closed-form solutions in the kernel regression paradigm, Ref. [36] replaces the neural network in the inner loop of the meta-learning framework with a kernel model. In the data matching framework, DC [37, 38] and DM [39, 40] aim to match the gradients and distribution of the distilled data with the original data during training. Trajectory matching [41] generates synthetic data by matching training trajectories.
Recently, the integration of dataset distillation into FL has attracted widespread attention. In DynaFed[42], the server generates distilled data by matching the trajectory after each aggregation and further updates the global model based on these data. Ref. [43] improves communication efficiency and test accuracy by constructing distilled data on each client to locally match the loss landscape of the original data. However, these works do not consider the use of D2D (device-to-device) networks to compensate for client heterogeneity. Ref. [44] introduces a method for extracting meta-knowledge from clients and exchanging it among clients, but this knowledge sharing may be less effective in scenarios with severe data heterogeneity.
In comparison with the aforementioned works, we design a new framework of HFL to address the issue of non-IID data. It extends the idea of constructing heterogeneous topological structure in DFL to HFL, and fully leverages the potential of heterogeneity among clients.
@@ III-A Dataset Distillation
The main idea of dataset distillation is to transfer a large dataset to a small dataset, while preserving the essential characteristics of the original dataset. Suppose that the original data set and the distilled dataset are, respectively, denoted by $\mathcal{T}=\left\{\left(\mathrm{x}_{t,i},y_{t,i}\right)\right\}_{i=1}^{|\mathcal{T}|}$ and $\mathcal{S}=\left\{\left(\mathrm{x}_{s,j},y_{s,j}\right)\right\}_{j=1}^{|\mathcal{S}|}$, where $\mathrm{x}_{t,i}\in\mathbb{R}^{d}$ and $y_{t,i}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $i$-th sample in the original dataset. $\mathrm{x}_{s,j}\in\mathbb{R}^{d}$ and $y_{s,j}\in\mathbb{R}^{n_{c}}$ are, respectively, the input data and the corresponding label of the $j$-th sample in the distilled dataset. $|\mathcal{T}|$ and $|\mathcal{S}|$ represent the sizes of the original dataset and the distilled dataset, respectively. Generally, the size of $\mathcal{T}$ is considerably larger than $\mathcal{S}$, i.e., $|\mathcal{S}|\ll|\mathcal{T}|$.
To realize dataset distillation, one of the most widely used methods is Kernel Inducing Points (KIP) [45, 36], which is a first-order meta-learning algorithm based on kernel ridge regression (KRR). Specifically, the KIP method is to find a synthetic dataset $\mathcal{S}$ that is approximate to the target dataset $\mathcal{T}$ by minimizing the following loss function
@@
where $\lambda>0$ is a regularization parameter. $y_{t}\in\mathbb{R}^{|\mathcal{T}|\times n_{c}}$ and $y_{s}\in\mathbb{R}^{|\mathcal{S}|\times n_{c}}$ are the labels of $\mathcal{T}$ and $\mathcal{S}$, respectively. ${K}_{X_{s}X_{s}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ and $K_{X_{t}X_{s}}\in\mathbb{R}^{|\mathcal{T}|\times|\mathcal{S}|}$ are two kernel matrices defined as $(K_{X_{s}X_{s}})_{ij}=k\left(\mathrm{x}_{s,i},\mathrm{x}_{s,j}\right)$ and $(K_{X_{t}X_{s}})_{ij}=k\left(\mathrm{x}_{t,i},\mathrm{x}_{s,j}\right)$, where $k(\cdot)$ denotes the kernel function. KIP is highly efficient for the distillation of datasets because it involves only first-order optimization.
@@ III-B Federated Learning
Consider a classic FL scenario with $N$ participants collaboratively training a global model. Each participant possesses a local dataset, represented as $\mathcal{D}_{i}=\left\{\left(X_{i},y_{i}\right)\right\},i=1,\cdots,N$, where $X_{i}\in\mathbb{R}^{n_{i}\times d}$ and $y_{i}\in\mathbb{R}^{n_{i}\times n_{c}}$ denote the tuple of input data points and their corresponding labels. $n=\sum_{i=1}^{N}{n_{i}}$, where $n_{i}=|\mathcal{D}_{i}|$ is the number of samples in $\mathcal{D}_{i}$. The objective of collaborative learning is to solve the following distributed parallel optimization problem
@@
In the process of training, each participating client trains the local model based on their local dataset and then transmits the local models to a server for global model aggregation. Specifically, in round $t$, the client $i$ trains its local model $\omega_{i}^{t}$ by minimizing the local loss function $f_{i}\left(\omega\right)$ using the gradient descent method. Subsequently, the server updates the global model $\omega^{t}$ through aggregation rules, e.g., FedAvg. If client data are IID, the above method can provide the optimal solution of (2). while if the client data are non-IID, the obtained global model may be drifted from the optimum. The communication cost of the above process comes from the local model upload and the global model download. By referring to Ref. [29], the communication cost of the widely used FedAvg is
@@
$a_{h}$ cluster head in the $h$-th heterogeneous cluster
$\mathcal{A}$ set of cluster head
$\mathcal{D}_{h}$ whole dataset on the $h$-th cluster head
@@
${\omega}_{h}^{t}$ local model on the $h$-th cluster head at communication round $t$
@@
$E_{1}$ number of local update steps in pre-train at label knowledge collection stage
$E_{2}$ number of local update steps in local training at model training stage
@@ III-C Problem Statement
In this paper, we consider a typical scenario of distributed learning with non-IID data, i.e., label distribution skew [46]. Our focus lies on addressing the issue of quantity-based label imbalance, where each client owns data with a fixed number of labels. Supposing $N_{c}$ denotes the total number of label classes, each client owns data with $C$ different labels. Data labels among different clients could overlap. If $C=N_{c}$, the data labels among all clients are the same, which can be regarded as quantity-based label balance. If $C<N_{c}$, the data labels among all clients are imbalanced. If $C=1$, it represents the worst-case scenario of quantity-based label imbalance, where the local dataset on each client contains only a single label. The quantity-based label imbalance has been verified to greatly decrease the performance of model training [14, 47].
Instead of traditional FL, we will design a new HFL framework that integrates dataset distillation to create approximately label-balanced local datasets, thereby alleviating global model drift caused by non-IID data. Furthermore, it should be noted that clients can communicate with each other.